OpenStudy (kkutie7):

A runner runs in a straight line with the average velocity of 5i-hat m/s for 5 minutes and then with an average velocity of 4i-hat m/s 4 minutes. a) what is her total displacement: b)What is her average velocity during this time c) suppose her goal is to run 5 km in 15 minutes. What average velocity must she have over the remaining portion of her run?

OpenStudy (kkutie7):

a) 5min*60s=300s 5m/s*300s=1500m 4min*60s=240s 4m/s*240s=960m 1500+960=2460 i-hatm

OpenStudy (kkutie7):

what do I do for b?

OpenStudy (kkutie7):

so displacement over what?

OpenStudy (tylerd):

so took her 9 min to run 2460. 6 minutes and 2.54km to go

OpenStudy (kkutie7):

what is her displacement?

OpenStudy (kkutie7):

@satellite73 @Hero